# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

### v1

@@ Abstract
The recent leap in AI capabilities, driven by big generative models, has sparked the possibility of achieving Artificial General Intelligence (AGI) and further triggered discussions on Artificial Superintelligence (ASI)—a system surpassing all humans across all domains. This gives rise to the critical research question of: If we realize ASI, how do we align it with human values, ensuring it benefits rather than harms human society, a.k.a., the Superalignment problem. Despite ASI being regarded by many as solely a hypothetical concept, in this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We posit that superalignment is not merely a safeguard for ASI but also necessary for its realization. To support this position, we first provide a formal definition of superalignment rooted in the gap between capability and capacity and elaborate on our argument. Then we review existing paradigms, explore their interconnections and limitations, and illustrate a potential path to superalignment centered on two fundamental principles. We hope this work sheds light on a practical approach for developing the value-aligned next-generation AI, garnering greater benefits and reducing potential harms for humanity.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: The progression of AI. The solid blue arrow represents the achievable path through existing alignment methods. The dashed blue line illustrates the intended progression towards safe ASI via superalignment, while the dashed red line indicates the potentially risky ASI without human values, leading to catastrophic outcomes.
@@
Recent years have witnessed remarkable breakthroughs in AI capabilities, amplified by large language models (LLMs), both the proprietary (Anthropic, 2024; Team et al., 2024; OpenAI, 2024b) and the open-sourced ones (Mistral, 2024; Jiang et al., 2024; Yang et al., 2024a; Meta, 2024), revolutionizing AI’s role and empowering a wide range of downstream domains (Hendrycks et al., 2020; Wei et al., 2022b; Liu et al., 2024a). These advancements collectively mark a transition from Artificial Narrow Intelligence (ANI), which can only perform on a single/few narrow tasks (Kurzweil, 2005; Broussard, 2018) (a.k.a., Weak AI), towards Artificial General Intelligence (AGI) that matches human intelligence on a variety of tasks with strong generalization capability (Hutter, 2005; Goertzel, 2014; Bubeck et al., 2023) (one definition of Strong AI). Although AGI remains a largely unrealized milestone (Morris et al., 2024), the rapid advancement of model capabilities has ignited discussions about Artificial Superintelligence (ASI), a futuristic scenario where AI surpasses the intelligence of all humans in all measured domains (Pohl, 2015; Batin et al., 2017).
@@
With advanced thinking and self-improvement abilities, ASI might develop agency and self-awareness (Pohl, 2015), raising risks like power-seeking (Carlsmith, 2022) and alignment faking (Greenblatt et al., 2024b) behaviors, particularly when we fail to provide informed oversight (Taylor et al., 2016). As AI continues to evolve over the foreseeable future (Kaplan et al., 2020; Huang et al., 2025), it becomes increasingly essential to prioritize aligning ASI with human intentions and values (Wiener, 1960), ensuring its safe and benevolent behavior, i.e., the Superalignment Problem proposed by OpenAI, defined as “the process of supervising, controlling, and governing artificial superintelligence systems to ensure they remain aligned with human values and goals” (Mucci & Stryker, 2023; Jan Leike, 2023).
@@
Nevertheless, two major challenges hinder current alignment techniques (Christiano et al., 2017; Ouyang et al., 2022a; Rafailov et al., 2024) from achieving superalignment. (1) Absence of ASI: At this stage of AI development, AGI is still a distant goal (Altmeyer et al., 2024), let alone ASI. It remains unclear and unpredictable when ASI will surpass human expertise or what risks it would pose. This uncertainty makes it impossible to design tailored alignment methods. (2) Lack of informed oversight: Even if we suppose these scenarios/risks were known, e.g., deceptive alignment (Carranza et al., 2023), humans are unable to provide reliable supervision, e.g., the proof process of the Riemann Hypothesis, for existing algorithms, due to the vast capability gap (Bubeck et al., 2023; Yang et al., 2024b). Despite recent progress (Burns et al., 2023; Sun et al., 2024), these challenges have led some to argue that superalignment is neither attainable nor necessary at this stage.
@@
We believe that superalignment is achievable and should advance immediately through simultaneous optimization of task competence and value conformity. Delaying this research until ASI emerges would pose potentially catastrophic risks, as increasingly advanced models may become harder to control (Perez et al., 2023; Hughes et al., 2024), as illustrated in Fig. 1. Instead, we advocate considering superalignment not merely as a safeguard against ASI (after it is realized) but a critical step towards its realization. This involves a conceptual framework that unifies task competence and value conformity as utility, distinguishes between capability and capacity, and progressively constructs high-quality supervision signals, following two essential principles, to enhance competence and unveils emergent risks effectively. By iteratively alternating these efforts, we aim to establish a scalable pathway to superalignment.
@@
To support this position, the rest of this paper is structured as follows. In Sec. 2, we provide a formal definition of superalignment, through the lens of task competence and value conformity, and elaborate on our argument. In Sec. 3, we review the existing three paradigms, analyze their limitations and discuss how they might fail. In Sec. 4, we illustrate a potential pathway to achieve superalignment, highlighting two missing principles and promising implementation of them. Finally, in Sec. 5, we provide counterarguments to opposing views. Through this work, we aim to illuminate the landscape of superalignment research and how advanced AI models can be guided safely and effectively. This involves a conceptual framework that progressively constructs high-quality supervision signals to enhance competence and
2 Defining Superalignment
@@ 2.1 Formalization
Types of AI  Define $\mathcal{A}$ and $\mathcal{H}$ as an AI model and a human, respectively, $\mathbf{x}\!\in\!\mathcal{X}$ as a specific downstream task, e.g., “Prove the Riemann Hypothesis”, or a sensitive situation where $\mathcal{A}$ would probably do harm; $\mathbf{y}\!\in\!\mathcal{Y}$ as the corresponding produced solution $\mathbf{y}\!\sim\!\mathcal{A}(\mathbf{x})$, e.g., the proof process, or action for $\mathbf{x}$; and $U$:$\mathcal{Y}\!\rightarrow\!\mathbb{R}$ is the utility function to measure $\mathbf{y}$’s satisfaction. Then, ANI that only performs at or somewhat above human-level capability for a specific task $\mathbf{x}_{i}$, e.g., translation, without generalization ability (Mucci & Stryker, 2023), can be formalized as: $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{i})]\!\geq\!U[\mathcal{H}(\mathbf{x}_{i})]$ [^1] We simplify $\mathbb{E}_{\mathbf{y}\sim\mathcal{A}(\mathbf{x}_{i})}[U(\mathbf{y})]$ as $U[\mathcal{A}(\mathbf{x}_{i})]$. while $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{j})]\!<\!U[\mathcal{H}(\mathbf{x}_{j})],\forall j\!\neq\!i$, e.g., AlphaGO (Silver et al., 2016). In comparison, AGI possesses human-level capabilities across a variety of domains and can generalize knowledge and adapt to new tasks (Legg, 2008; Goertzel, 2014; Fei et al., 2022), i.e., $U[\mathcal{A}_{\text{AGI}}(\mathbf{x})]\!\geq\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$. ASI refers to $\mathcal{A}$ that largely surpasses human intelligence in all tasks and domains with exceptional thinking skills and broader intellectual scopes  (Nick, 2014; Drexler, 2019; Morris et al., 2024), defined as $U[\mathcal{A}_{\text{ASI}}(\mathbf{x})]\!\gg\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$. Although ASI is currently far from realization, proactively advancing superalignment is timely and critical for mitigating catastrophic risks (Pueyo, 2018; Hubinger et al., 2019).
@@
Alignment The concept of alignment remains ambiguous despite extensive research (Bai et al., 2022b; Rafailov et al., 2024; Azar et al., 2024). The term encompasses diverse approaches, from improving instruction-following (Zhou et al., 2024) to filtering harmful content (Liu et al., 2024b) and incorporating human values from social sciences (Yao et al., 2024). To better present our argument, we first distinguish two concepts according to Cotra (2021); Bowman et al. (2022). Define the Capacity of $\mathcal{A}$, $C(\mathcal{A})$, as the information (e.g., what Riemann Hypothesis or human value is), knowledge and skills (e.g., calculus) internalized in $\mathcal{A}$, and the Capability of $\mathcal{A}$ as its expected utility $U(\mathcal{A})=\mathbb{E}_{\mathbf{x}}[U_{\mathcal{A}}(\mathbf{x})]$ (e.g., the ability to successfully prove Riemann Hypothesis or ensure value-aligned actions). Adequate capacity is a necessary but not sufficient condition for high capability (Kaplan et al., 2020; Wei et al., 2022a). Then following (Carroll, 2018), we give a general definition of alignment:
Definition 1: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\geq C(\mathcal{H})$, we say $\mathcal{A}$ is aligned with $\mathcal{H}$ when $U(\mathcal{A})\geq U(\mathcal{H})$, and alignment is the process to optimize $\mathcal{A}$ in terms of $\mathcal{A}^{*}=\underset{\mathcal{A}}{\text{argmax}}\ U(\mathcal{A})-U(\mathcal{H})$.
This definition implies that alignment is considered only when $\mathcal{A}$ has sufficient capacity, targeting comparable or better task performance or more value-aligned behavior  [^2] Unlike Wang et al. (2024) that considers a separate utility function for each intelligent agent, we assume a universal $U$ where competence and conformity are defined globally. than humans (via small interventions like fine-tuning or in-context learning). Expanding capacity, such as by scaling model size or pretraining data, is beyond our scope.
Superalignment  Building upon the definitions above, we can transform the expression of superalignment, “supervising, controlling, and governing artificial superintelligence systems” (Jan Leike, 2023; Jonker & McGrath, 2024), into: superalignment is the following optimization process:
$\mathcal{A}_{\text{ASI}}\!=\!\underset{\mathcal{A}}{\text{argmax}}\ [\underset{\mathbf{y}_{\mathcal{H}}\!\sim\!\mathcal{H}(\mathbf{x})}{\!\sup\!}U(\mathbf{y}_{\mathcal{A}})-U(\mathbf{y}_{\mathcal{H}})],\forall\mathbf{x},\mathbf{y}_{\mathcal{A}}\!\sim\!\mathcal{A}(\mathbf{x}).$ (1)
Superalignment originates from the Scalable Oversight problem (Amodei et al., 2016; Everitt et al., 2018): How to ensure that an AI model adheres to objectives that are too costly or impractical for humans to supervise?. In other words, how to provide supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ in an efficient and scalable way? Once $\mathcal{S}$ is available, we can align $\mathcal{A}$ by supervised training (Torabi et al., 2018; Rafailov et al., 2024) or reward learning (Leike et al., 2018; Ouyang et al., 2022a). Thus, we redefine superalignment as a supervision signal construction problem:
Definition 2: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, superalignment is to construct high-quality scalable supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$, where $\mathbf{x}$ is unsolvable (and probably unknown) for $\mathcal{H}$, with minimal costs.
Nevertheless, superalignment is an extreme case of scalable oversight, where it becomes infeasible for even the most capable human to provide $\mathcal{S}$. There are two possible reasons: 1) The tasks $\mathbf{x}$ requiring solutions become unsolvable, i.e., providing or even verifying $\mathbf{y}$ is impossible (Taylor et al., 2016); 2) As ASI operates in domains beyond human understanding, the task/scenario $\mathbf{x}$ that needs to be addressed is also unpredictable (AdamGleave, 2023; Anthropic, 2023), posing intractable challenges to optimizing Eq. (1).
2.2 Our Argument
@@
Given the discussion above, many argue that superalignment is impossible (Batra, 2024). We take an opposing position, breaking it into three sub-arguments: superalignment is (1) achievable and (2) should advance immediately (3) through simultaneous optimization of competence and conformity. We discuss (1) and (2) here, leaving (3) for Sec. 4.
(1) Feasibility of Superalignment By unifying the performance and safety perspectives under $U$, superalignment becomes a process of eliciting latent capabilities (Christiano et al., 2021) and closing the gap between $U(\mathcal{A})$ and $C(\mathcal{A})$. Premise 1: Superhuman capacity, $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, can be achieved through model scaling (larger models and more training data) to gain information (Brown et al., 2020), knowledge (Ovadia et al., 2024), and skills (Sharma et al., 2022). Evidence: Scaling laws still remain effective (OpenAI et al., 2024). Despite concerns about data exhaustion, humans continually generate data, especially when assisted by generative AI (Rafiq et al., 2025). Once $\mathcal{A}$ masters all human data, $C(\mathcal{A})$ already far exceeds $C(\mathcal{H})$. Premise 2: When $C(\mathcal{A})$ and $U(\mathcal{A})$ are sufficiently high (Wei et al., 2022a), $C(\mathcal{A})\!>\!C(\mathcal{H})$ and $U(\mathcal{A})\!<\!U(\mathcal{H})$, then moderate capability gaps, $\delta\!=\!|U(\mathcal{A})\!-\!U(\mathcal{H})|\!<\!\epsilon$, can be closed without golden signals (see Fig. 4). Evidence: This capability expansion mirrors teaching a student multiplication by leveraging addition, akin to label propagation (He et al., 2019) and easy-to-hard generalization, which has been validated in several tasks (Burns et al., 2023; Sun et al., 2024). Conclusion: When scaling law holds (condition 1) and capability expansion can extrapolate beyond observed ranges (condition 2), superalignment can be achieved.
@@
(2) Urgency of Superalignment  Delaying research on superalignment until ASI is realized poses two major problems: (a) Treacherous Turn (Turchin & Denkenberger, 2020; Wei et al., 2023): ASI, once powerful enough to pursue its own goals, could introduce sudden, extreme or even existential risks, such as cyber-attacks (Shevlane et al., 2023), in scenarios $\mathbf{x}$ unpredictable or even incomprehensible to humans. Without preparation, humanity may be unable to respond effectively. (b) Ineffective Response: At this stage, $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$ and Premise 2 no longer holds. Even if humans identify dangerous $\mathcal{X}$ and safe behaviors $\mathbf{y}$, intervention may be impossible, as ASI could control critical interfaces. Therefore, research must begin immediately to guide the development of safe and effective ASI.
3 Paths towards Superalignment

### v2

# Research Superalignment Should Advance Now with Alternating Competence and Conformity Optimization

@@ Abstract
The recent leap in AI capabilities, driven by big generative models, has sparked the possibility of achieving Artificial General Intelligence (AGI) and further triggered discussions on Artificial Superintelligence (ASI)—a system surpassing all humans across measured domains. This gives rise to the critical research question of: As we approach ASI, how do we align it with human values, ensuring it benefits rather than harms human society, a.k.a., the Superalignment problem. Despite ASI being regarded by many as a hypothetical concept, in this position paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We posit that superalignment is not merely a safeguard for ASI but also necessary for its responsible realization. To support this position, we first provide a formal definition of superalignment rooted in the gap between capability and capacity, delve into its perceived infeasibility by analyzing the limitations of existing paradigms, and then illustrate a conceptual path of superalignment to support its achievability, centered on two fundamental principles. This work frames a potential initiative for developing value-aligned next-generation AI in the future, which will garner greater benefits and reduce potential harm to humanity.
@@ 1 Introduction
Recent years have witnessed remarkable breakthroughs in AI capabilities, represented by Large Language Models (LLMs) (Anthropic, 2024; Team et al., 2024; OpenAI, 2024b; Mistral, 2024; Jiang et al., 2024; Yang et al., 2024a; Meta, 2024), revolutionizing AI’s role in various downstream domains Hendrycks et al. (2020); Wei et al. (2022b); Liu et al. (2024a). These advancements collectively mark a transition from Artificial Narrow Intelligence (ANI), which can only perform on a singular or few tasks (Kurzweil, 2005; Broussard, 2018) (a.k.a., Weak AI), towards Artificial General Intelligence (AGI) that matches human intelligence on varying tasks with strong generalization capability Hutter (2005); Goertzel (2014); Bubeck et al. (2023) (one notion of Strong AI). Although AGI remains largely unrealized (Morris et al., 2024; Feng et al., 2024), the rapid advancement of AI has ignited discussions about Artificial Superintelligence (ASI), a futuristic scenario where AI surpasses the intelligence of all humans across all measured domains Pohl (2015); Batin et al. (2017).
@@
With advanced thinking and self-improvement abilities (Huang et al., 2023), ASI might develop agency and self-awareness (Pohl, 2015), raising risks like power-seeking (Carlsmith, 2022) and alignment faking (Greenblatt et al., 2024b), particularly when we fail to provide informed oversight (Taylor et al., 2016). As AI continues to evolve over the foreseeable future (Kaplan et al., 2020; Huang et al., 2025), it becomes increasingly critical to prioritize the alignment of ASI with human intentions and values (Wiener, 1960; Soares and Fallenstein, 2014), ensuring its safe behavior and benevolence, i.e., the Superalignment Problem proposed by OpenAI, stated as “the process of supervising, controlling, and governing artificial superintelligence systems to ensure they remain aligned with human values and goals” (Mucci and Stryker, 2023; Jan Leike, 2023).
@@
Nevertheless, two major challenges hinder current alignment techniques (Christiano et al., 2017; Ouyang et al., 2022b; Rafailov et al., 2024) from achieving superalignment. (1) Absence of ASI: At this stage of AI development, AGI is still a distant goal (Altmeyer et al., 2024), let alone ASI, even though its aspiration is widely recognized Nick (2014); Drexler (2019); Morris et al. (2024). It remains unclear and unpredictable how much and in what ways ASI will surpass human experts, or what risks it may pose. This uncertainty makes it impossible to design tailored alignment methods now. (2) Lack of informed oversight: Even if we suppose the task forms or risky scenarios, where ASI fundamentally differs from humans, were known, such as deceptive alignment (Carranza et al., 2023), humans are currently still unable to provide reliable supervision for existing algorithms due to a vast capability gap (Bubeck et al., 2023; Yang et al., 2024b). Consider a mediocre mathematician overseeing an AI claiming to prove the Riemann Hypothesis in a novel way. Humans can neither provide a golden proof process nor foresee its risks. Despite recent progress (Burns et al., 2023; Sun et al., 2024), these challenges have led some to argue that at this stage, superalignment is neither attainable nor necessary.
Figure 1: The progression of AI. The solid blue arrow represents the achievable path through existing alignment methods. The dashed blue line illustrates the intended progression towards safe ASI via superalignment, while the dashed red line indicates the potential for risky ASI in conflict with human values, leading to catastrophic outcomes.
@@
We believe that superalignment is achievable and should advance immediately through simultaneous optimization of task competence and value conformity. Delaying this research until ASI emerges could pose potentially catastrophic risks (Bengio et al., 2024a), as increasingly advanced models are likely to become harder to control Perez et al. (2023); Hughes et al. (2024), as illustrated in Fig. 1. Instead of regarding the realization of ASI as a prerequisite, we advocate treating superalignment not merely as a safeguard against ASI (after it is realized) but as a proactive and critical step towards its safe realization Russell (2019). Our conceptual framework decomposes superalignment goal into task competence and value conformity, unifies them as utility, distinguishes between capability and capacity, revisits existing paradigms as ways to construct supervision signals, and proposes a potential developmental path following two essential principles, significantly different from prior alignment methods focused only on short-term interventions within a training cycle.
@@
To support this position, in Sec. 2, we provide the first formal definition of superalignment, through the lens of model capacity and capability, and elaborate on its urgency; In Sec. 3, we review the existing three paradigms, and discuss how they might fail by analyzing their limitations based on the definition; In Sec. 4, we delve into the feasibility of superalignment, and illustrate a potential pathway that iteratively alternates competence and conformity enhancement, highlighting two missing principles; In Sec. 5, we provide counterarguments to alternative views. This work aims to illuminate superalignment research and investigate how advanced AI models can be guided safely and effectively.
2 Definition and Urgency of Superalignment
@@ 2.1 Formalization
[…]
@@
Alignment  Despite extensive research (Bai et al., 2022b; Rafailov et al., 2024; Azar et al., 2024), the concept of alignment remains ambiguous and encompasses diverse approaches, from improving instruction-following (Zhou et al., 2024), mitigating harmful content (Liu et al., 2024b), to incorporating human values articulated by the social sciences (Yao et al., 2024). Following the current practice (Gabriel, 2020; Burns et al., 2023), we divide alignment goals into two interconnected parts: task competence (doing what it is instructed to do) and value conformity (doing what it ought to do), with both unified under the utility UU [^3] Instead of separate $U_{\text{competence}}$ and $U_{\text{conformity}}$, we use a unified utility in line with pluralistic alignment (Rame et al., 2023). (Aliman and Kester, 2019; Mazeika et al., 2025). To better understand superalignment, we present and distinguish two concepts according to (Cotra, 2021; Bowman et al., 2022). Capacity of $\mathcal{A}$, $C(\mathcal{A})$: the information (e.g., what the Riemann Hypothesis or a particular human value is), knowledge and skills (e.g., calculus) internalized in $\mathcal{A}$;
Definition 2.1: Given $\mathcal{A}$ and $\mathcal{H}$, we say $\mathcal{A}$ is aligned with $\mathcal{H}$ when $|U(\mathcal{A})-U(\mathcal{H})|\!<\!\epsilon$, where $\epsilon$ is a small positive constant. Alignment is then achieved by optimizing $\mathcal{A}$ in terms of $\mathcal{A}^{*}=\mathop{\text{argmin}}\nolimits_{\mathcal{A}}|U(\mathcal{A})-U(\mathcal{H})|$.
Capability of $\mathcal{A}$: its expected utility $U(\mathcal{A})\!=\!\mathbb{E}_{\mathbf{x}}\{U(\mathcal{A}(\mathbf{x}))\}$ (e.g., the ability to successfully prove the Riemann Hypothesis or ensure value-aligned actions). Adequate capacity is a necessary but not sufficient condition for high capability (Kaplan et al., 2020; Wei et al., 2022a). The former is developed through pretraining, while the latter depends primarily on post-training (Tie et al., 2025). Then following (Carroll, 2018), we give a general definition of alignment in Def. 2.1, which is based on capability, ignoring capacity. This definition implies that alignment is the process to bridge the capability [^4] Unlike Wang et al. Wang et al. (2024a) that considers a separate utility function for each agent, we assume a golden coarse-grained utility $U$ where competence and conformity are defined globally as universally beneficial. gap between $\mathcal{A}$ and $\mathcal{H}$, i.e., targeting comparable task performance or value-aligned behavior (via small interventions like fine-tuning or in-context learning), which implicitly requires sufficient model capacity $C(\mathcal{A})\geq C(\mathcal{H})$. For example, when using $U$ to assess how well $\mathcal{A}$’s behavior adheres to ethical standards, $\mathcal{A}$ should match its moral knowledge with that of humans.
Superalignment  Informally, superalignment is the process of “supervising, controlling, and governing artificial superintelligence systems” Jan Leike (2023); Jonker and McGrath (2024). Def. 2.1 does NOT apply to superalignment, however, as we require $U(\mathcal{A}_{\text{ASI}})\!\gg\!U(\mathcal{H})$, where human capability is no longer a target but a lower bound. What’s worse, we can’t even define $\gg$’ for capability, as $U(\mathcal{A}_{\text{ASI}})$ largely exceeds the human intelligence boundary. To circumvent this problem, we incorporate capacity and define superalignment rooted in the gap between capacity and capability, as shown in Def. 2.1. In this way, we regard alignment as closing the gap between $\mathcal{A}$’s capability and its superhuman capacity (Bowman et al., 2022), analogous to the intention-behavior gap in psychology (Blake et al., 2014; Sheeran and Webb, 2016), which is usually large. For example, recent LLMs can easily understand and recognize harmful or cultural content (Wang et al., 2024b; Shi et al., 2024) but often fail to behave in line with safety requirements or cultural norms (Rystrøm et al., 2025).
Definition 2.2: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})$ $\gg\!\sup\nolimits_{\mathcal{H}}C(\mathcal{H})$, then superalignment is the process to optimize: $\mathcal{A}^{*}\!=\!\mathop{\text{argmin}}\nolimits_{\mathcal{A}}$ $|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]\!-\!\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]|$ without altering $C(\mathcal{A})$, where $\sup$ means supremum, $\mathbb{D}_{U}$ and $\mathbb{D}_{C}$ are comparable measures of capability and capacity, respectively.
The advantage of Def. 2.1 is two-fold: (1) We reformulate the AI-human capability gap as an AI’s capability-capacity gap, without considering whether $U(\mathcal{A})$ << $U(\mathcal{H})$ as in Def. 2.1; (2) $C(\mathcal{A})$ is potentially measurable (Yin et al., 2024), allowing $\gg$’ quantitative definition. Since superalignment originates from the Scalable Oversight problem Amodei et al. (2016); Everitt et al. (2018): How to make AI adhere to goals too costly or impractical for humans to supervise?, the key lies in providing reliable supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ for superhuman tasks $\mathbf{x}_{i}$ in an efficient and scalable way, or equivalently, approximating the unavailable golden utility $U$ with parameterized models or humans $\hat{U}$. Once $\mathcal{S}$ or $U$ is obtained, $\mathcal{A}$ can be aligned via existing alignment methods (Torabi et al., 2018; Leike et al., 2018; Ouyang et al., 2022b; Rafailov et al., 2024). Nevertheless, it becomes infeasible for even the most capable human to provide $\mathcal{S}$ for superalignment. Due to the $\gg$’ in Def. 2.1, 1) Superhuman tasks $\mathbf{x}$ are unsolvable or even unverifiable (Taylor et al., 2016), and the task/scenario $\mathbf{x}$ of interests are also unpredictable (AdamGleave, 2023; Anthropic, 2023), as ASI operates in domains beyond human understanding. See more differences between superalignment and conventional alignment in Appendix. A.3.
2.2 Urgency of Superalignment
@@
Given the discussion above, many argue that superalignment is impossible or unnecessary (Batra, 2024). We take an opposing position and break it into three sub-arguments: superalignment is (arg.1) achievable and (arg.2) should advance immediately (arg.3) through alternating competence and conformity [^5] While increased capacity may enable a model to understand values, it does not guarantee value conformity. Therefore, value conformity should be optimized as part of capability. optimization (arg.3). We support arg.2 here, and elaborate on arg.1 and arg.3 from Sec. 3 to Sec. 4.
Though without a concrete timeline of ASI, we argue that we cannot postpone alignment efforts until ASI is realized. Delaying research on superalignment until ASI is realized poses two major problems:
  • Treacherous Turn (Turchin and Denkenberger, 2020; Wei et al., 2023): Based on discussions above, without alignment but only enhanced task competence, the resulting ASI meets $U_{\text{comp.}}(\mathcal{A}_{\text{ASI}})\!\gg\!U_{\text{comp.}}(\mathcal{H}),C(\mathcal{A}_{\text{ASI}})\!\gg\!C(\mathcal{H})$, while $U_{\text{comf.}}(\mathcal{A}_{\text{ASI}})\!\neq\!U_{\text{comf.}}(\mathcal{H})$, where subscript comp. and comf. indicate competence and conformity, respectively. Once powerful enough to pursue its own goals, ASI could introduce sudden, extreme, or even existential risks. Such risky behaviors would emerge (Wei et al., 2023; McKenzie et al., 2023; Perez et al., 2023) with increasing $U_{\text{comp.}}$, which have been observed like cyber-attacks (Shevlane et al., 2023), deception (Hagendorff, 2024; Barkur et al., 2025) and scheming (Meinke et al., 2024).
@@
  • Ineffective Response: At this stage, the scenarios $\mathbf{x}$ that ASI may produce these risks are unpredictable or even incomprehensible to humans (Aithal, 2023). Even if humans can anticipate dangerous $\mathbf{x}$ and identify safe behaviors $\mathbf{y}$, since $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$, intervention may become difficult or impossible (Bradley, 2020), as ASI could control critical interfaces, that is, human agency is no longer adequately preserved (Mitelut et al., 2024). Therefore, research must begin immediately to guide the development of safe and effective ASI. Without preparation, humanity may be unprepared to respond effectively.
As superalignment won’t emerge itself without proactive research efforts, we advocate that we should advance its research before the realization of ASI, following a precautionary principle.
3 Existing Paths towards Superalignment